OpenStudy (anonymous):

Trig: A point is rotating with uniform circular motion on a circle of radius r. Find w (omega) if r = 6 cm and v = 3 cm/sec Help me set it up please

OpenStudy (anonymous):

r = 6cm v = 3cm/sec it takes you 4 seconds for 1 rotation 4 seconds = 1 period w = 2 * Pi / Periodlength w = 2* Pi / 4 = Pi /2 = 1.57

OpenStudy (anonymous):

How do you know it takes you 4 seconds for 1 rotation?

OpenStudy (anonymous):

r = 6cm , the diameter is r * 2 = 12 12 /v = 4

OpenStudy (anonymous):

by the way the answer is 0.5 rad/sec that is why I'm kind of confuse

OpenStudy (anonymous):

yeah its the same haha. r = 6, v = 3cm/sec v = 0.5rad/sec

OpenStudy (anonymous):

btw. v = 0.5rad/sec is not the answer . you want to find omega.

OpenStudy (anonymous):

that is what it says on the book